OpenStudy (anonymous):

Evaluate the definite integral [e^(1/x^(5))/x^(6)]dx from 1 to 2

OpenStudy (anonymous):

the answer i got was \[e-\sqrt[32]{e}\] but it says its wrong. can someone help me figure out where i went wrong.

OpenStudy (sidsiddhartha):

take 1/x^5=u the -5/x^6*dx=du now just substitute

OpenStudy (anonymous):

\[I=\int\limits_{1}^{2}\frac{ e ^{\frac{ 1 }{ x^5 }} }{ x^6 }dx\] \[put~\frac{ 1 }{ x^5}=t,~x ^{-5}=t,-5 x ^{-6}dx=dt,\frac{ dx }{ x^6 }=\frac{ -1 }{ 5 }dt\] when x=1,t=1,when x=2,t=1/32 \[I=-\frac{ 1 }{ 5 }\int\limits_{1}^{\frac{ 1 }{ 32 }} e^t dt=?\]

OpenStudy (anonymous):

thanks both of you it worked. apparently i'm an airhead

OpenStudy (sidsiddhartha):

lol

OpenStudy (anonymous):

yw